# Impact of Weak Lensing Mass Mapping Algorithms on Cosmology Inference

### v1

@@ 4.5 Emulator
To obtain a prediction for our HOS for any given point in the parameter space $\left[\Omega_{\textrm{m}},\sigma_{8},h,w_{0}\right]$ within the range of our simulations, we construct an emulator using the Gaussian Process Regression (GPR) method 222 We use the GPyTorch (version 1.12) Python package (Gardner et al., 2018) to implement the GPR.. For each bin of the HOS histograms, we train a separate GPR model on all realizations of the HOS from the Cosmology Training Set. We found that training the GPR models on the full set of realizations, rather than just the mean HOS histograms, improved the emulator’s performance. This approach enables the GPR models to predict the scaled HOS values in each histogram bin as a function of the cosmological parameters while simultaneously capturing the variability across realizations through a Gaussian likelihood.
@@
The GPR training involves optimizing the model hyperparameters by maximizing the marginal log-likelihood using the Adam optimizer. We employ a Radial Basis Function kernel with Automatic Relevance Determination, allowing the model to learn different length scales for each cosmological parameter. To ensure stability and efficiency, the HOS values are rescaled by their mean value in each bin during training, and the scaling factors are recorded for subsequent use in predictions.
@@
We then validate the emulator by performing a leave-one-out cross-validation, where we train the GPR on all but one of the 25 cosmologies in the Cosmology Training Set, and then predict the HOS for the left-out cosmology. We repeat this process for all the cosmologies in the Cosmology Training Set, and we calculate the percentage error between the predicted and the true HOS. Because the training is done on one less cosmology than the full set, the leave-one-out cross-validation provides an upper limit on the error of the emulator. In particular, the error is over-estimated in the case that the left-out cosmology is on the edge of the training volume, since in this case the emulator has to extrapolate from the other training points, instead of interpolating. We find that overall the emulator has percentage errors of less than $5\%$ for all the HOS, which is acceptable for our purposes.
@@
However, as discussed in Harnois-Deraps et al. (2024), the small number of available cosmologies in the cosmo-SLICS simulations used for training the emulator can lead to significant biases in the posterior distributions of the cosmological parameters. To mitigate these biases arising from the Gaussian Process emulator altogether, following Davies et al. (2024), we used the emulator’s prediction of HOS at the given fiducial cosmology as our mock ”observed” data vector. This choice is also made for simplicity and presentation purposes, as it assures that the contours will be centred around the fiducial cosmology, making it easier to compare the constraining power of the HOS for the different mass mapping methods. To ensure that this choice does not significantly alter the results of the analysis, we also perform the inference analysis using a mock data vector calculated from a cosmo-SLICS simulation at the fiducial cosmology, but in a simpler setting where the GPR is more robust. The results of this analysis are presented in the Appendix B.
@@ 4.7 MCMC and posterior distributions
We perform a Markov Chain Monte Carlo (MCMC) analysis employing the emcee [^3] https://github.com/dfm/emcee Python package. We assume flat priors for the parameters $\Omega_{\textrm{m}}\in[0.1019,0.5482]$, $\sigma_{8}\in[0.4716,1.3428]$, $h\in[0.6034,0.8129]$, and $w_{0}\in[-1.9866,-0.5223]$, which are the boundaries of the training volume. We use the Gaussian likelihood function in Eq. (10), and a model-independent covariance, as described in Section 4.4. To verify the convergence of the chains, we use the Gelman-Rubin diagnostics (Gelman & Rubin, 1992).
@@ 5 Results
In this section we present the results of our inference analysis for the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$, using the pipeline described in the previous section. We show the parameter constraints, estimated from peak counts and wavelet peak counts, and we investigate the impact of the choice of the mass mapping method on these constraints.
5.1 Mono Scale Peak counts
Figure 6 shows the constraints obtained from single-scale peak counts for the KS, iKS, and MCALens methods. We observe a substantial difference in the constraints obtained from the different mass mapping methods, with the MCALens method providing the tightest constraints, followed by the KS and the iKS method.
We further quantify these results by calculating the Figure of Merit (FoM) for each pair of parameters, defined as:
@@ 5.1 Mono Scale Peak counts
Table 2: Figure of Merit (FoM) for each pair of parameters, as well as for the full set of parameters, from the single-scale peak counts, for the Kaiser-Squires (KS), iterative Kaiser-Squires (iKS), and MCALens methods.
Figure 6: Constraints on the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$ from the single-scale peak counts, for the KS (gray), iKS (red), and MCALens (blue) methods. The contours represent the $68\%$ and $95\%$ confidence levels, and the black dashed lines indicate the true parameter values of our fiducial ($\Lambda$CDM) cosmology.
@@ 5.2 Wavelet peak counts
Figure 7 shows the constraints obtained from the wavelet peak counts, for the KS, iKS, and MCALens methods. As in the case of the single-scale peak counts, we observe a clear difference in the constraints obtained from the different mass mapping methods, with the MCALens method delivering the tightest constraints, followed by the KS and the iKS method. Table LABEL:tab:mspc_fom shows the FoM for each pair of parameters, as well as for the full set of parameters, for the KS, iKS, and MCALens methods.
@@
To assess the contribution of different wavelet scales across various methods, we also perform the inference analysis incrementally by including an increasing number of scales. We begin by obtaining constraints using the $[16^{\prime},32^{\prime},\textrm{coarse}]$ scales. Next, we successively add finer scales to this set, starting with the $[8^{\prime}]$ scale, followed by the $[4^{\prime}]$ scale, and, finally the $[2^{\prime}]$ scale.
The results are shown in Figs. 8 and 9 for the KS and MCALens methods, respectively. In Fig. 8, we observe that KS constraints improve significantly when including the $[8^{\prime}]$ scale, while the $[4^{\prime}]$ and $[2^{\prime}]$ scales do not provide significant additional information. In contrast, Fig 9 shows that the MCALens constraints improve significantly with the inclusion of every smaller scale. This result highlights that the MCALens method is able to extract more information from the smaller scales of the convergence field, and that is where it gains most of its constraining power over the KS and iKS methods.
Figure 7: Constraints on the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$ from the wavelet peak counts with 6 wavelet scales $[2^{\prime},4^{\prime},8^{\prime},16^{\prime},32^{\prime},\textrm{coarse}]$, for the KS (gray), iKS (red), and MCALens (blue) methods. The contour lines are described in Fig. 6.
@@
Table 3: Figure of Merit (FoM) for each pair of parameters, as well as for the full set of parameters, from the wavelet peak counts, for the Kaiser-Squires (KS), iterative Kaiser-Squires (iKS), and MCALens methods.
Figure 8: Constraints on the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$ from the wavelet peak counts using KS inversion and including different number of wavelet scales. The green contours were obtained with three scales: $[16^{\prime},32^{\prime},\textrm{coarse}]$, the gray contours were obtained by adding a fourth scale at $8^{\prime}$. The red contours were obtained by adding a fifth scale at $4^{\prime}$, and the blue contours include a sixth scale at $2^{\prime}$. The contour lines are described in Fig. 6.
Figure 9: Constraints on the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$ from the wavelet peak counts using MCALens inversion including different number of wavelet scales. The green contours were obtained with three scales: $[16^{\prime},32^{\prime},\textrm{coarse}]$, and the gray contours were obtained by adding a fourth scale at $8^{\prime}$. The red contours were obtained by adding a fifth scale at $4^{\prime}$, and the blue contours include a sixth scale at $2^{\prime}$. The contour lines are described in Fig. 6.
@@ 6 Conclusions
In this study, we investigated the impact of various mass mapping algorithms on cosmological inference using HOS derived from weak-lensing data. We have used a set of $25$ simulated cosmologies to train an emulator for the HOS, and we have used a set of $1,240$ realizations of the HOS for a fixed cosmology to estimate the covariance matrices. We have then performed a Bayesian analysis with an MCMC method to estimate the posterior distributions of the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$, using peak counts and wavelet peak counts as our data vectors. Our analysis shows that the choice of the mass mapping method has a significant impact on the constraints on the cosmological parameters, with the MCALens method providing the tightest constraints. The improvement by using MCALens is $\sim 17\%$ in the FoM compared to the KS method for the mono-scale analysis, and $\sim 157\%$ for the multi-scale analysis. We also find no significant difference in the contours created with the iKS and the KS methods, indicating that forward modelling the inversion is sufficient to account for edge effects for both peak counts and multi-scale peak count statistics. Finally, we confirmed previous results that wavelet peak counts provide tighter constraints than single-scale peak counts. Constraints from MCALens increase the most by using a multi-scale approach because it captures additional information from smaller scales over KS.
These results highlight that a careful choice of mass mapping method to accurately reconstruct the convergence field can significantly improve constraints on the cosmological parameters from HOS.
@@ Acknowledgements.
This work was supported by the TITAN ERA Chair project (contract no. 101086741) within the Horizon Europe Framework Program of the European Commission, and the Agence Nationale de la Recherche (ANR-22-CE31-0014-01 TOSCA). LB is supported by the PRIN 2022 project EMC2 - Euclid Mission Cluster Cosmology: unlock the full cosmological utility of the Euclid photometric cluster catalog (code no. J53D23001620006). This work used the CANDIDE computer system at the IAP supported by grants from the PNCG, CNES, and the DIM-ACAV and maintained by S. Rouberol. The authors thank Nicolas Martinet, Virginia Ajani and Joachim Harnois-Deraps for useful discussions.
Appendix A Mass mapping algorithms
A.1 Kaiser-Squires (KS)
@@
where $\mathbf{A}$ is the convolution matrix (which corresponds to the linear transformation from ideal $\kappa$-field to ideal $\gamma$-field), and $\bm{n}$ is the noise vector. The matrix $\mathbf{A}$ can be decomposed in Fourier space as $\mathbf{A}=\mathbf{F}\mathbf{P}\mathbf{F}^{*}$, where $\mathbf{F}$ is the discrete Fourier transform matrix, and $\mathbf{P}$ is the Fourier-space filter that relates the pure shear signal (in the absence of noise) to the convergence
@@
Here, $\tilde{\bm{\gamma}}=\tilde{\gamma}_{1}+i\tilde{\gamma}_{2}$ and $\tilde{\bm{\kappa}}=\tilde{\kappa}_{E}+i\tilde{\kappa}_{B}$ are the complex representations of the two shear components and the convergence $E$ and $B$-modes, in Fourier space, respectively.
The original method to solve the mass-mapping problem was proposed by Kaiser & Squires (1993).
This method directly inverts the relation between the observed shear $\gamma$ and the convergence $\kappa$ in Fourier space, as given by Eq. (2). In practice, this is done by directly applying the Fourier-space filter $\mathbf{P}^{-1}$ to the shear field:
@@
While the KS method is the simplest and computationally fastest mass-mapping technique, it does not account for the noise, missing data, or the boundaries of the survey. This makes it suboptimal in the application to real data, leading to significant noise amplification, edge effects, and internal artifacts in the reconstructed maps, along with leakage between the $E$ and $B$-modes of the convergence field. Typically, to reduce the noise in the reconstructed maps, some Gaussian smoothing is applied to the map. This, however, also leads to a loss of information at small scales and suppression of peaks.
To rigorously account for the noise covariance matrix ${\operatorname*{\bm{\Sigma}}}_{n}$ and the mask M, we need to minimize the following Eq.:
@@
which can be solved using the following iteration:
@@
where $\mu=\min(\operatorname*{\bm{\Sigma_{n}}})$. This iterative algorithm generalizes the standard KS solution, to take into account the noise covariance matrix and the mask. More details can be found in Starck et al. (2021), Appendix B.1.
A.2 Iterative Kaiser-Squires with DCT Inpainting (iKS)
Inpainting methods are techniques used to fill in missing data in an image or signal by estimating the missing values from known data. Sparse inpainting based on the Discrete Cosine Transform was proposed in Starck et al. (2005a, b) and used in the context of mass-mapping by Pires et al. (2009) to fill in areas of missing data in the convergence field.
Inpainting can be very easily included in Eq. 15, so that one now minimizes the following equation:

### v2

@@ 4.5 Emulator
To obtain a prediction for our HOS for any given point in the parameter space $\left[\Omega_{\textrm{m}},\sigma_{8},h,w_{0}\right]$ within the range of our simulations, we constructed an emulator using the Gaussian process regression (GPR) method 333 We used the GPyTorch (version 1.12) Python package (Gardner et al. 2018) to implement the GPR.. For each bin of the HOS histograms, we trained a separate GPR model on all realisations of the HOS from the cosmology training set. We find that training the GPR models on the full set of realisations, rather than just the mean HOS histograms, improved the emulator’s performance. This approach enables the GPR models to predict the scaled HOS values in each histogram bin as a function of the cosmological parameters while simultaneously capturing the variability across realisations through a Gaussian likelihood.
@@
The GPR training involves optimising the model hyperparameters by maximising the marginal log-likelihood using the Adam optimiser. We employed a radial basis function kernel with automatic relevance determination, allowing the model to learn different length scales for each cosmological parameter. To ensure stability and efficiency, the HOS values are rescaled by their mean value in each bin during training, and the scaling factors are recorded for subsequent use in predictions.
@@
We then validated the emulator by performing a leave-one-out cross-validation, where we trained the GPR on all but one of the 25 cosmologies in the cosmology training set, and then predicted the HOS for the left-out cosmology. We repeated this process for all the cosmologies in the cosmology training set, and we calculated the percentage error between the predicted and the true HOS. Because the training is done on one fewer cosmology than the full set, the leave-one-out cross-validation provides an upper limit on the error of the emulator. In particular, the error is over-estimated in the case that the left-out cosmology is on the edge of the training volume, since in this case the emulator has to extrapolate from the other training points, instead of interpolating. We find that overall the emulator has percentage errors of less than $5\%$ for all the HOS, which is acceptable for our purposes.
@@
However, as discussed in Harnois-Déraps et al. (2024), the small number of available cosmologies in the cosmo-SLICS simulations used for training the emulator can lead to significant biases in the posterior distributions of the cosmological parameters. To mitigate these biases arising from the Gaussian process emulator altogether, following Davies et al. (2024), we used the emulator’s prediction of HOS at the given fiducial cosmology as our mock ‘observed’ data vector. This choice is also made for simplicity and presentation purposes, as it assures that the contours will be centred around the fiducial cosmology, making it easier to compare the constraining power of the HOS for the different mass-mapping methods. To ensure that this choice does not significantly alter the results of the analysis, we also perform the inference analysis using a mock data vector calculated from a cosmo-SLICS simulation at the fiducial cosmology, but in a simpler setting where the GPR is more robust. The results of this analysis are presented in Appendix B.
@@ 4.7 MCMC and posterior distributions
We performed a Markov chain Monte Carlo (MCMC) analysis employing the emcee [^4] https://github.com/dfm/emcee Python package. We assumed flat priors for the parameters $\Omega_{\textrm{m}}\in[0.1019,0.5482]$, $\sigma_{8}\in[0.4716,1.3428]$, $h\in[0.6034,0.8129]$, and $w_{0}\in[-1.9866,-0.5223]$, which are the boundaries of the training volume. We used the Gaussian likelihood function in Eq. (10) and a model-independent covariance, as described in Sect. 4.4. To verify the convergence of the chains, we used the Gelman-Rubin diagnostics (Gelman & Rubin 1992).
@@ 5 Results
In this section we present the results of our inference analysis for the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$, using the pipeline described in the previous section. We show the parameter constraints, estimated from peak counts and wavelet peak counts, and we investigate the impact of the choice of the mass-mapping method on these constraints.
5.1 Mono-scale peak counts
Figure 6 shows the constraints obtained from single-scale peak counts for the KS, iKS, and MCALens methods. We observe a substantial difference in the constraints obtained from the different mass-mapping methods, with the MCALens method providing the tightest constraints, followed by the KS and the iKS method.
We further quantified these results by calculating the figure of merit (FoM) for each pair of parameters, defined as
@@ 5.1 Mono-scale peak counts
where $\tilde{F}$ is the marginalised Fisher matrix, estimated as the inverse of the covariance matrix among the parameters of interest, calculated from our MCMC chains, and $n$ is the number of parameters. The FoM is a measure of the constraining power of the data on the parameters, with higher values indicating tighter constraints. Table 2 shows the FoM for each subset of parameters, as well as for the full set of parameters, for the KS, iKS, and MCALens methods.
@@
Table 2: FoMs for parameters from single-scale peak counts, for the KS, iKS, and MCALens methods
@@
Figure 6: Constraints on the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$ from the single-scale peak counts, for the KS (grey), iKS (red), and MCALens (blue) methods. The contours represent the $68\%$ and $95\%$ confidence levels, and the dashed black lines indicate the true parameter values of our fiducial ($\Lambda$CDM) cosmology.
@@ 5.2 Wavelet peak counts
Figure 7 shows the constraints obtained from the wavelet peak counts, for the KS, iKS, and MCALens methods. As in the case of the single-scale peak counts, we observe a clear difference in the constraints obtained from the different mass-mapping methods, with the MCALens method delivering the tightest constraints, followed by the KS and the iKS method. Notably, KS and iKS produce very similar contours, indicating again that the inpainting applied by iKS has only a minimal effect for peak counts. Table 3 shows the FoM for each pair of parameters, as well as for the full set of parameters, for the KS, iKS, and MCALens methods.
@@
To assess the contribution of different wavelet scales across various methods, we also performed the inference analysis incrementally by including an increasing number of scales. We began by obtaining constraints using the $[16^{\prime},32^{\prime},\textrm{coarse}]$ scales. Next, we successively added finer scales to this set, starting with the $8^{\prime}$ scale, followed by the $4^{\prime}$ scale, and, finally the $2^{\prime}$ scale.
The results are shown in Figs. 8 and 9 for the KS and MCALens methods, respectively. In Fig. 8, we observe that KS constraints improve significantly when including the $8^{\prime}$ scale, while the $4^{\prime}$ and $2^{\prime}$ scales do not provide significant additional information. In contrast, Fig 9 shows that the MCALens constraints improve significantly with the inclusion of every smaller scale. This result highlights that the MCALens method is able to extract more information from the smaller scales of the convergence field, and that is where it gains most of its constraining power over the KS and iKS methods.
Figure 7: Constraints on the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$ from the wavelet peak counts with six wavelet scales $[2^{\prime},4^{\prime},8^{\prime},16^{\prime},32^{\prime},\textrm{coarse}]$, for the KS (grey), iKS (red), and MCALens (blue) methods. The contour lines are described in Fig. 6.
Table 3: FoMs for parameters from wavelet peak counts, for the KS, iKS, and MCALens methods
@@
Figure 8: Constraints on the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$ from the wavelet peak counts using KS inversion and including different numbers of wavelet scales. The green contours were obtained with three scales: $[16^{\prime},32^{\prime},\textrm{coarse}]$. The grey contours were obtained by adding a fourth scale at $8^{\prime}$. The red contours were obtained by adding a fifth scale at $4^{\prime}$, and the blue contours include a sixth scale at $2^{\prime}$. The contour lines are described in Fig. 6.
Figure 9: Constraints on the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$ from the wavelet peak counts using MCALens inversion and including different numbers of wavelet scales. The green contours were obtained with three scales: $[16^{\prime},32^{\prime},\textrm{coarse}]$. The grey contours were obtained by adding a fourth scale at $8^{\prime}$, the red contours a fifth scale at $4^{\prime}$, and the blue contours a sixth scale at $2^{\prime}$. The contour lines are described in Fig. 6.
@@ 6 Conclusions
We investigated the impact of various mass-mapping algorithms on cosmological inference using HOS derived from weak-lensing data. We used a set of $25$ simulated cosmologies to train an emulator for the HOS, and we used a set of $1,240$ realisations of the HOS for a fixed cosmology to estimate the covariance matrices. We then performed a Bayesian analysis with an MCMC method to estimate the posterior distributions of the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$, using peak counts and wavelet peak counts as our data vectors.
Our analysis shows that the choice of the mass-mapping method has a significant impact on the constraints on the cosmological parameters, with the MCALens method providing the tightest constraints. The improvement gained by using MCALens is $\sim 17\%$ in the FoM compared to the KS method for the mono-scale analysis, and $\sim 157\%$ for the multi-scale analysis. We find no significant difference in the contours created with the iKS and KS methods, indicating that forward-modelling the inversion is sufficient to account for edge effects for both peak counts and multi-scale peak count statistics. However, we note that this conclusion does not necessarily generalise to all HOS. Certain HOS, such as the bispectrum, are much more sensitive to masked regions because they rely on information across Fourier modes that can be significantly impacted by missing data. In such cases, a dedicated study would be required to determine whether an inpainting approach or another treatment of the masked area is necessary to ensure reliable results. Finally, we confirm previous results, that wavelet peak counts provide tighter constraints than single-scale peak counts. Constraints from MCALens are tightest when using a multi-scale approach because it captures additional information from smaller scales compared to KS.
These results highlight that a careful choice of mass-mapping method when attempting to accurately reconstruct the convergence field can significantly improve constraints on the cosmological parameters from HOS.
@@ Acknowledgements.
This work was supported by the TITAN ERA Chair project (contract no. 101086741) within the Horizon Europe Framework Program of the European Commission, and the Agence Nationale de la Recherche (ANR-22-CE31-0014-01 TOSCA). LB is supported by the PRIN 2022 project EMC2 - Euclid Mission Cluster Cosmology: unlock the full cosmological utility of the Euclid photometric cluster catalog (code no. J53D23001620006). This work used the CANDIDE computer system at the IAP supported by grants from the PNCG, CNES, and the DIM-ACAV and maintained by S. Rouberol. The authors thank Nicolas Martinet, Virginia Ajani and Joachim Harnois-Deraps for useful discussions. The authors also thank Joachim Harnois-Deraps for providing the simulations used in this work.
Appendix A Mass-mapping algorithms
A.1 Kaiser-Squires
@@
where $\mathbf{A}$ is the convolution matrix (which corresponds to the linear transformation from ideal $\kappa$-field to ideal $\gamma$-field), and $\bm{n}$ is the noise vector. The matrix $\mathbf{A}$ can be decomposed in Fourier space as $\mathbf{A}=\mathbf{F}^{*}\mathbf{P}\mathbf{F}$, where $\mathbf{F}$ is the discrete Fourier transform matrix, and $\mathbf{P}$ is the Fourier-space filter that relates the pure shear signal (in the absence of noise) to the convergence
@@
Here, $\tilde{\bm{\gamma}}=\tilde{\gamma}_{1}+i\tilde{\gamma}_{2}$ and $\tilde{\bm{\kappa}}=\tilde{\kappa}_{E}+i\tilde{\kappa}_{B}$ are the complex representations of the two shear components and the convergence $E$ and $B$ modes, in Fourier space, respectively.
The original method for solving the mass-mapping problem was proposed by Kaiser & Squires (1993). This method directly inverts the relation between the observed shear $\gamma$ and the convergence $\kappa$ in Fourier space, as given by Eq. (2). In practice, this is done by directly applying the Fourier-space filter $\mathbf{P}^{-1}$ to the shear field:
@@
This is the relation that was used for the standard KS reconstruction in our analysis.
While the KS method is the simplest and computationally fastest mass-mapping technique, it does not account for the noise, missing data, or the boundaries of the survey. This makes it suboptimal in the application to real data, leading to significant noise amplification, edge effects, and internal artefacts in the reconstructed maps, along with leakage between the $E$ and $B$ modes of the convergence field. Typically, to reduce the noise in the reconstructed maps, some Gaussian smoothing is applied to the map. This, however, also leads to a loss of information at small scales and suppression of peaks.
To rigorously account for the noise covariance matrix ${\operatorname*{\bm{\Sigma}}}_{n}$ and the mask M, we needed to minimise the following equation:
@@
This can be solved using the following iterative process:
@@
where $\mu=\min(\operatorname*{\bm{\Sigma_{n}}})$. This iterative algorithm generalises the standard KS solution, to take into account the noise covariance matrix and the mask. For the special case of a diagonal covariance matrix with all diagonal elements equal to 1 and no mask (i.e. $\textbf{M}=1$) this iterative algorithm converges to the standard KS solution. In particular, if we set $\mu=1/2$, the update simplifies to the KS inversion in one iteration. More details can be found in Starck et al. (2021), Appendix B.1.
A.2 Iterative Kaiser-Squires with DCT inpainting
Inpainting methods are techniques used to fill in missing data in an image or signal by estimating the missing values from known data. Sparse inpainting based on the discrete cosine transform (DCT) was proposed in Starck et al. (2005a, b) and used in the context of mass-mapping by Pires et al. (2009) to fill in areas of missing data in the convergence field.
Inpainting can be very easily included in Eq. 15, so that one now minimises the following equation: